Given that a 50 foot tall amazon women from a science fiction movie (bad sci-fi) is similar to a normal sized woman, determine her weight given that the average woman is 5.5 ft tall and weighs 140 lbs.
Answer by Theo(13342)

the amazon woman has the same proportions as the normal woman.

the normal woman is 5.5 feet tall and weight 140 pounds.

if you let x equal the weight of the amazon woman, then the amazon woman is 50 feet tall and x pounds.

you get a ratio that says that 5.5 / 140 = 50 / x.

cross multiply to get 5.5 * x = 140 * 50

divide both sides of this equation by 5.5 to get x = (140 * 50) / 5.5.

solve for x to get x = 1272.727272... pounds.

that's equivalent to 1272 and 8/11 pounds.

if you look at the height to weight ratio, the normal wouman is 5.5 / 140 = .0392857143 and the amazon woman is 50 / (1272 + 8/11) = the same.

they're in proportion.